# S-GRPO: Early Exit via Reinforcement Learning in Reasoning Models

### v1

@@ Abstract
As Test-Time Scaling emerges as an active research focus in the large language model community, advanced post-training methods increasingly emphasize extending chain-of-thought (CoT) generation length, thereby enhancing reasoning capabilities to approach Deepseek R1-like reasoning models. However, recent studies reveal that reasoning models (even Qwen3) consistently exhibit excessive thought redundancy in CoT generation. This overthinking problem stems from conventional outcome-reward reinforcement learning’s systematic neglect in regulating intermediate reasoning steps. This paper proposes Serial-Group Decaying-Reward Policy Optimization (namely S-GRPO), a novel reinforcement learning method that empowers models with the capability to determine the sufficiency of reasoning steps, subsequently triggering early exit of CoT generation. Specifically, unlike GRPO, which samples multiple possible completions (parallel group) in parallel, we select multiple temporal positions in the generation of one CoT to allow the model to exit thinking and instead generate answers (serial group), respectively. For the correct answers in a serial group, we assign rewards that decay according to positions, with lower rewards towards the later ones, thereby reinforcing the model’s behavior to generate higher-quality answers at earlier phases with earlier exits of thinking. Empirical evaluations demonstrate compatibility with state-of-the-art reasoning models, including Qwen3 and Deepseek-distill models, achieving 35.4$\sim$61.1sequence length reduction with 0.72$\sim$6.08accuracy improvements across GSM8K, AIME 2024, AMC 2023, MATH-500, and GPQA Diamond benchmarks.
@@ 1 Introduction
Test-Time Scaling snell2024scalingllmtesttimecompute demonstrates a strong correlation between extended chain-of-thought (CoT) and enhanced reasoning capabilities in Large Language Models, which is more effective than scaling model parameters kaplan2020scalinglawsneurallanguage . The success of DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability and GPT-o1 openai2025learning has further sparked the research interest in reasoning models xu2025largereasoningmodelssurvey within the LLM community. Furthermore, reinforcement learning shao2024deepseekmathpushinglimitsmathematical ; bai2022training ; ouyang2022training ; schulman2017proximal ; ramesh2024group (RL) in post-training has demonstrated significant potential in stimulating long chain-of-thought generation and strengthening deep thinking.
@@
However, recent studies chen2025think23overthinkingo1like ; team2025kimi have uncovered a critical limitation in reasoning models: their inherent tendency toward redundant thought processes, namely Overthinking chen2025think23overthinkingo1like ; cuadron2025danger . They frequently generate unnecessarily lengthy reasoning sequences wu2025more ; cuadron2025danger that incorporate extraneous information and non-essential reasoning steps. This redundant thought not only inflates computational overhead but also reduces reasoning accuracy through deviation from valid reasoning pathways to incorrect conclusions yang2025dynamicearlyexitreasoning . We attribute this phenomenon to the inherent nature of 0/1 outcome-reward RL (exemplified by GRPO shao2024deepseekmathpushinglimitsmathematical ), where training with only final outcome rewards is unable to determine when reasoning steps have reached sufficiency.
@@
As shown in Figure 1 (left), the standard outcome-reward RL, like GRPO, tends to sample a query multiple times in parallel to obtain a parallel group. All reasoning chains and corresponding conclusions within the parallel group receive 0/1 outcome rewards, which reinforce correct ones regardless of overthinking in intermediate steps. While this approach successfully aggregates pass@k reasoning capability into pass@1, the neglect of regulating intermediate reasoning results in inefficient inference. Conversely, as illustrated in Figure 1 (right), we constructe a serial group by sequentially collecting multiple completions during single CoT generation, and prioritize rewards for earlier correct ones, aiming to enable models to acquire and exert the complete reasoning capability in the initial generation phases, triggering early exit before overthinking occurs.
@@
Motivated by this, we propose Serial-Group Decaying-Reward Policy Optimization (S-GRPO), a simple yet effective modification to address standard outcome-reward RL (GRPO)’s inability to regulate intermediate reasoning processes. Specifically, during S-GRPO training, we construct a serial group for each query through a two-phase rollout: the initial rollout generates a complete reasoning path, followed by subsequent rollouts that perform early-exit interventions at different positions to generate intermediate answers, which are collected with their corresponding truncated reasoning paths into the serial group. On this basis, we assign the rewards that decay according to their order of early exit for the correct ones within the serial group. S-GRPO culminates in computing serial-group relative advantages and using their policy gradient to update model parameters.
@@
  • The proposed S-GRPO algorithm empowers models to generate higher-quality reasoning paths at early stages of CoTs while implementing early exit when the sufficiency of generated thought is met – a strategy particularly suitable as final-stage optimization in post-training pipelines.
  • Extensive experiments across MATH-500, AMC, GSM8K, AIME and GPQA benchmarks with Qwen3 and Deepseek-series reasoning models, demonstrate 0.74$\sim$6.08accuracy improvement alongside 35.4$\sim$61.1average token reduction, establishing an efficiency-accuracy synergistic improvement.
@@ 2 Method
The proposed Serial-Group Decaying-Reward Policy Optimization (S-GRPO) is a novel reinforcement learning mechanism that innovatively leverages rule-based outcome rewards to regulate intermediate reasoning processes. During the training of S-GRPO, we exit early and generate answers at different positions in a complete CoT, thus constructing a serial group and assigning rewards that decrease according to the position, which reinforces the model to generate high-quality thoughts earlier and exit early when the generated thought is sufficient. In particular, the framework is divided into three stages: Serial-Group Generation, Decaying Reward Policy, and Advantage Computation and Parameter Update.
@@ 2.1 Serial-Group Generation
GRPO is originally designed to enable reasoning models to achieve their pass@k potentiality under pass@1 settings. To accomplish this, GRPO performs parallel sampling of multiple CoTs per query to construct a parallel group, then rewards the correct ones. In contrast, S-GRPO aims to achieve efficient reasoning through sufficient-reasoning early exit. Specifically, we perform early-exit interventions at different positions within a single CoT to construct a serial group, thereby allowing the RL training can compare the thought sufficiency between reasoning paths of different lengths. It consists of two stages: Full Thought Rollout and Early-exit Thought Rollout.
@@ 2.1.1 Full Thought Rollout
In the Full Thought Rollout, the model generates a complete reasoning path step by step $\{O^{0}=T_{1},T_{2},\ldots,T_{n},</think>,C_{0}\}$ for each query. To simulate diverse reasoning sufficiency scenarios and train the model to handle early exits effectively, we adopt random-length truncation during training. Specifically, the reasoning sequence is randomly truncated at $m$ different temporal positions $P_{i}$, where $P_{i}=T_{i}\quad and\quad i\sim\text{Uniform}(1,n).$ This randomness ensures that the model learns to evaluate sufficiency implicitly across varying reasoning depths and avoids overfitting to fixed-length reasoning paths.
@@ 2.1.2 Early-exit Thought Rollout
In the Early-exit Thought Rollout, the policy model extracts several early-exit thougths $\{CoT^{1},CoT^{2},...,CoT^{m}\}=\{O^{0}[:P_{1}],O^{0}[:P_{2}]...,,O^{0}[:P_{m}]\}$ of different lengths between the first token and multiple trancated-postion tokens of the full thought, respectively, and generates answers $C_{1},C_{2},\ldots,C_{m}$ for all early-exit CoTs.
@@
2.2 Decaying Reward Policy
To encourage models to produce adequate and correct reasoning steps at earlier stages of CoT generation for accurate problem-solving, we propose a Decaying Reward Policy. This mechanism assigns rewards based on the correctness of answers generated from the two-time rollouts ( $C_{1},C_{2},\ldots,C_{m},and\ C_{0}$ ), while decaying rewards according to their order of early exits. For the answer $C^{i}$ of each response $O^{i}$, the reward $r^{i}$ is defined as follows:
$r^{i}=\begin{cases}\frac{1}{2^{N_{\text{right}}-1}},&\text{if }C^{i}\text{ is correct,}\\ 0,&\text{if }C^{i}\text{ is incorrect.}\end{cases}$
Where $N_{\text{right}}$ refers to the accumulated number of correct answers up to and including the current position. The Decaying Reward Policy is designed with dual objectives: (1) Exponential decay for correct answers: We employ exponentially diminishing rewards to prioritize the quality of earlier reasoning steps, which is overlooked by binary 0/1 outcome rewards. (2) Zero reward for incorrect answers: This ensures correctness-first optimization, preventing degradation of the model’s core reasoning capabilities.
The design of the above two stages balances the trade-off between reasoning sufficiency and efficiency, guiding the model toward producing accurate and concise reasoning sequences.
@@ 2.3 Advantage Computation and Parameter Update
After obtaining the decaying rewards, GRPO calculates the advantage for each sample based on the serial group’s rewards. Specifically, the mean of the rewards within the group is computed, and the advantage for each sample is calculated using the formula: $\hat{A}_{i}=r_{i}-\text{mean}(r_{i})$. Here, for training stability, the standard deviation is removed from the advantage computation compared to the GRPO. Subsequently, the computed advantage for each sample is broadcast to all corresponding response tokens. Finally, parameter updates are performed based on the advantage values of each sample (Algorithm 1). The optimization objective is as follows:
@@
5:Sample $m$ temporal positions $\{P_{1},P_{2},\ldots,P_{m}\}$ where $P_{i}=T_{i}$ and $i\sim\text{Uniform}(1,n)$
@@
8:for each position $P_{i}$ where $i\in\{1,2,\ldots,m\}$ do
9:     Append $T_{\text{</think>}}$ after position $P_{i}$ to form $CoT^{i}=(T_{1},T_{2},\ldots,T_{i},T_{\text{</think>}})$
@@ 3.1 Experimental Setup
To comprehensively assess the models’ capabilities across a range of reasoning tasks, we select five popular benchmarks that reflect diverse levels of difficulty: GSM8K (cobbe2021trainingverifierssolvemathgsm8k, ) is a well-curated collection of 1,319 problems in elementary mathematics. This benchmark is specifically designed to evaluate multi-step reasoning in foundational math tasks. Problems typically require between two and eight sequential operations, relying primarily on basic arithmetic performed over multiple intermediate steps. AIME 2024 (aime, ) includes 30 advanced problems selected from the 2024 edition of the American Invitational Mathematics Examination. This highly regarded competition tests participants’ ability to reason mathematically across a broad range of topics, including arithmetic, algebra, combinatorics, geometry, number theory, and probability—core components of secondary-level mathematics. AMC 2023 (AMC2023, ) consists of 40 questions spanning key areas such as algebra, geometry, number theory, and combinatorics. As part of the American Mathematics Competitions organized by the Mathematical Association of America (MAA), AMC aims to foster problem-solving skills and identify promising young mathematicians. MATH-500 math500hendrycks2021measuringmathematicalproblemsolving presents a set of challenging problems drawn from high school-level mathematical competitions. To ensure comparability with prior studies, we use the same subset of 500 problems originally compiled by OpenAI for evaluation purposes (lightman2023letsverifystepstep, ). In addition to these mathematical evaluations, we also examine performance on scientific reasoning tasks, using the following benchmark: GPQA (rein2023gpqagraduatelevelgoogleproofqa, ) is a rigorously constructed dataset containing graduate-level questions in physics, chemistry, and biology. Notably, even domain experts with PhDs achieve only 69.7% accuracy on this benchmark o1 . For our experiments, we focus on the highest-quality subset, known as GPQA Diamond, which contains 198 carefully chosen questions.
@@
We compare S-GRPO with various existing efficient reasoning methods, including the training-free, output-based approach DEER yang2025dynamicearlyexitreasoning , the off-policy preference optimization method ConCISE qiao2025conciseconfidenceguidedcompressionstepbystep , and on-policy RL-based approaches such as original GRPO, RL + Length Penalty arora2025traininglanguagemodelsreason , and ShorterBetter yi2025shorterbetterguidingreasoningmodels . DEER makes early-exit decisions during the inference phase, guided by the confidence scores of intermediate answers. ConCISE generates more concise chains of thought by integrating specialized prompt tokens and applying early-exit mechanisms during inference, followed by preference optimization using SimPO to further encourage succinct reasoning. RL + Length Penalty assigns reward values based on the deviation of each individual response length from the mean, penalizing longer responses. ShorterBetter promotes shorter yet accurate reasoning paths by assigning higher rewards to compact chains that yield correct answers based on GRPO.
@@
S-GRPO aims to enhance the correctness performance while reducing inference length, thereby enabling more efficient reasoning. To this end, we selected Accuracy (i.e., pass@1) and Token Number (Tokens) as the evaluation metrics. Given the instability associated with generating long sequences in large reasoning models, as well as the limited number of samples in certain benchmark datasets, we conducted multiple runs for each evaluation and report the averaged results in the tables. Specifically, we ran 16 trials on AIME-2024 and AMC-2023, 8 trials on MATH-500 and GPQA Diamond, and 4 trials on GSM8K.
@@
During the S-GRPO, we use a learning rate of $1\times 10^{-6}$ and randomly choose 8 temporal positions for each query. As we use on-policy mode, the generation batch size and train batch size are the same and equal to 128*8. For the GRPO, we also use a learning rate of $1\times 10^{-6}$ and set the generation batch size and train batch size to 128*8. For the “RL + Length Penalty”, we refer to the settings in its paper arora2025traininglanguagemodelsreason and choose the scalar parameter ($\alpha$) of 0.2. For all experiments, Adam kingma2014adam is used as the standard optimizer.
@@
Method GSM8K AIME24 AMC23 MATH-500 GPQA Overall
@@
Vanilla 92.4 1833 55.4 13232 77.2 9693 85.8 5590 50.1 15385 70.02 9147
DEER 88.8 917 53.3 10971 87.5 4142 91.8 2431 47.5 5280 73.78+3.76 4748-48.1%
$\textit{ConCISE}_{SFT}$ 92.9 832 52.1 9751 92.0 2244 50.0 5892
$\textit{ConCISE}_{SimPO}$ 92.1 715 48.3 7745 91.0 1946 48.0 4859
GRPO 93.2 1767 55.0 13451 87.5 9887 93.6 5317 50.7 15817 76.00+5.98 9248+1.1%
RL + Length Penalty 92.4 1062 51.9 7464 86.9 3540 92.2 2451 49.1 3984 74.50+4.48 3700-59.5%
ShorterBetter 53.3 5288 75.9 2580
S-GRPO 93.8 906 56.0 7377 87.5 3494 92.4 2252 50.8 3751 76.10+6.08 3556-61.1%
@@
Vanilla 94.2 2129 64.4 11099 90.5 5527 93.5 3844 59.2 6034 80.36 5727
DEER 93.3 982 70.0 10335 90.0 4349 91.4 2753 57.1 4767 80.36+0.0 4637-19.0%
GRPO 95.3 2120 65.8 13504 91.9 6595 84.0 4471 58.9 7354 79.18-1.18 6809+18.9%
RL + Length Penalty 94.7 775 55.0 7950 88.1 3396 92.4 1993 56.0 4380 77.24-3.12 3699-34%
S-GRPO 96.2 724 64.4 6712 91.9 3352 93.6 2146 59.3 3334 81.08+0.72 3253-35.4%
@@
Vanilla 95.4 2370 74.1 15326 91.3 9452 93.4 5577 55.6 8741 81.90 8293
DEER 95.5 981 76.7 11287 95.0 6198 93.4 3208 52.5 3104 82.62+0.72 4956-40.2%
GRPO 95.8 2355 72.7 15154 92.8 8983 94.4 5440 55.8 8819 82.30+0.4 8150-1.7%
RL + Length Penalty 95.4 1323 73.8 9666 93.4 5042 94.2 3247 56.2 5293 82.60+0.7 4914-40.7%
S-GRPO 96.1 1292 77.3 8810 95.0 5962 95.2 3166 57.7 5271 84.26+2.36 4922-40.6%
@@
Vanilla 95.5 1909 75.4 14116 96.9 7576 95.2 5078 58.8 7576 84.36 7251
DEER 95.5 908 76.7 10333 95.0 5099 94.8 2987 57.1 2435 83.82-0.54 4352-40.0%
GRPO 96.1 1956 77.7 14544 98.4 8000 95.8 5140 59.3 7966 85.46+1.1 7521+3.7%
RL + Length Penalty 95.8 1090 74.8 9056 96.6 5059 95.8 2866 59.4 4949 84.46+0.1 4604-36.5%
S-GRPO 96.3 952 77.9 8932 97.8 4537 96.4 2652 60.6 4537 85.80+1.14 4322-40.4%

### v2

@@ Abstract
As Test-Time Scaling emerges as an active research focus in the large language model community, advanced post-training methods increasingly emphasize extending chain-of-thought (CoT) generation length, thereby enhancing reasoning capabilities to approach Deepseek R1-like reasoning models. However, recent studies reveal that reasoning models (even Qwen3) consistently exhibit excessive thought redundancy in CoT generation. This overthinking issue arises from the inherent limitations of conventional outcome-reward reinforcement learning, which systematically overlooks the regulation of intermediate reasoning processes. This paper introduces Serial-Group Decaying-Reward Policy Optimization (S-GRPO), a novel reinforcement learning paradigm that enables models to implicitly evaluate the sufficiency of intermediate reasoning steps, thereby facilitating early exit in CoT generation. Unlike GRPO, which samples multiple possible reasoning paths in parallel (parallel group), S-GRPO only samples one reasoning path and serially selects multiple temporal positions from the path to exit thinking and directly generate answers (serial group). For correct answers within a serial group, rewards gradually decrease based on the exit positions along the reasoning path from front to back. This design encourages the model to produce more accurate and concise thoughts, while also incentivizing early thinking termination when appropriate. Empirical evaluations demonstrate that S-GRPO is compatible with state-of-the-art reasoning models, including Qwen3 and Deepseek-distill. Across diverse benchmarks such as GSM8K, AIME 2024, AMC 2023, MATH-500, and GPQA Diamond, S-GRPO achieves a substantial reduction in sequence length (35.4$\sim$61.1%) while simultaneously improving accuracy (absolute 0.72$\sim$6.08%).
@@ 1 Introduction
Test-Time Scaling snell2024scalingllmtesttimecompute demonstrates a strong correlation between extended chain-of-thought (CoT) and enhanced reasoning capabilities in Large Language Models, which is more effective than scaling model parameters kaplan2020scalinglawsneurallanguage . The success of DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability and GPT-o1 openai2025learning has further sparked the research interest in reasoning models xu2025largereasoningmodelssurvey within the LLM community. Furthermore, reinforcement learning shao2024deepseekmathpushinglimitsmathematical ; bai2022training ; ouyang2022training ; schulman2017proximal ; ramesh2024group (RL) in post-training has demonstrated significant potential in stimulating long chain-of-thought generation and strengthening deep-thinking capabilities.
@@
However, recent studies chen2025think23overthinkingo1like ; team2025kimi have identified a critical limitation in reasoning models: their tendency to engage in redundant thought processes, a phenomenon referred to as Overthinking chen2025think23overthinkingo1like ; cuadron2025danger . They frequently generate unnecessarily lengthy reasoning sequences cuadron2025danger ; wu2025more , including irrelevant information and superfluous reasoning steps. This redundant thought inflates computational overhead and even undermines reasoning accuracy by diverting the model from valid reasoning pathways to incorrect ones yang2025dynamicearlyexitreasoning . We attribute this issue to the inherent limitations of 0/1 outcome-reward RL (e.g., GRPO shao2024deepseekmathpushinglimitsmathematical ), where reliance on final outcome rewards fails to identify when intermediate reasoning steps are sufficient.
@@
As shown in Figure 1 (left), a standard outcome-reward RL, such as GRPO, tends to sample a query multiple times in parallel to obtain a parallel group. All reasoning chains and corresponding conclusions within the parallel group receive 0/1 outcome rewards, which reinforces correct outcomes but overlooks the presence of overthinking or inefficiencies in intermediate reasoning steps. While this approach successfully aggregates pass@k reasoning capability into pass@1, the neglect of regulating intermediate reasoning results in inefficient inference. Conversely, as illustrated in Figure 1 (right), we construct a serial group by sequentially generating multiple completions during a single CoT process. By prioritizing rewards for earlier correct completions, this approach encourages models to demonstrate complete reasoning capabilities in the initial phases, enabling early exit and preventing overthinking.
@@
Motivated by this, we propose Serial-Group Decaying-Reward Policy Optimization (S-GRPO), a simple yet effective modification to address standard outcome-reward RL (GRPO)’s inability to regulate intermediate reasoning processes. During S-GRPO training, we construct a serial group for each query using a two-phase rollout process. In the first phase, a complete reasoning path is generated. In the second phase, subsequent rollouts introduce early-exit interventions at different positions along the reasoning path generated in the first phase, producing intermediate answers. Finally, the serial group is formed by combining the complete reasoning path from the first phase with the intermediate answers appended to its corresponding truncated reasoning path. On this basis, we assign the rewards that decay according to their order of early exit for the correct ones within the serial group. S-GRPO culminates in computing serial-group relative advantages and using their policy gradient to update model parameters.
@@
  • The proposed S-GRPO algorithm enables models to produce higher-quality reasoning paths during the early stages of CoT generation, and implicitly early-exit once sufficiency is achieved. S-GRPO preserves the integrity of the original reasoning process via a two-stage rollout procedure, ensuring that the model’s pre-existing reasoning abilities are not compromised, which is well-suited as the final stage of post-training.
  • Extensive experiments across GSM8K, AIME 2024, AMC 2023, MATH-500, and GPQA benchmarks with Qwen3 and Deepseek-series reasoning models, demonstrate 0.72$\sim$6.08absolute accuracy improvement alongside 35.4$\sim$61.1average token reduction, establishing an efficiency-accuracy synergistic improvement.
@@ 2 Method
The proposed Serial-Group Decaying-Reward Policy Optimization (S-GRPO) is a novel reinforcement learning mechanism that innovatively leverages rule-based outcome rewards to regulate intermediate reasoning processes and improve reasoning efficiency. During S-GRPO training, the LLM exits early and generates answers at different positions within a complete CoT, forming a serial group. Rewards for correct answers are assigned based on position, with earlier thinking exits receiving relatively higher rewards. This strategy encourages the model to produce high-quality reasoning earlier and terminate once sufficient reasoning is achieved. In particular, the framework is divided into three stages: Serial-Group Generation, Decaying Reward Strategy, and Advantage Computation and Parameter Update.
@@ 2.1 Serial-Group Generation
GRPO is originally designed to enable reasoning models to achieve their pass@k potential under pass@1 settings. To achieve this, GRPO generates multiple CoTs in parallel for each query, forming a parallel group, and rewards only those with correct answers. In contrast, S-GRPO aims to achieve efficient reasoning through sufficient-reasoning early exit. Specifically, we perform early-exit interventions at different positions within a single CoT to construct a serial group, thereby allowing the RL training to compare the thought sufficiency in different positions along the reasoning path. It consists of two stages: Full Thought Rollout and Early-exit Thought Rollout.
@@ 2.1.1 Full Thought Rollout
In the Full Thought Rollout stage, the model generates a complete reasoning path sequentially for each query, represented as $\{O^{0}=T_{1},T_{2},\ldots,T_{n},</think>,C_{0}\}$. To expose the model to diverse reasoning sufficiency scenarios and enhance its ability to handle early exits effectively, we adopt random-length truncation during training. Specifically, the reasoning sequence is truncated at $m$ randomly selected temporal positions $P_{i}$, where $P_{i}=T_{i}\quad and\quad i\sim\text{Uniform}(1,n)$, ensuring that the truncation points are uniformly distributed across the reasoning path. This randomness allows the model to implicitly evaluate reasoning sufficiency across varying depths and prevents overfitting to specific reasoning path lengths.
@@ 2.1.2 Early-exit Thought Rollout
In the Early-exit Thought Rollout stage, the policy model extracts several early-exit reasoning paths $\{CoT^{1},CoT^{2},...,CoT^{m}\}=\{O^{0}[:P_{1}],O^{0}[:P_{2}]...,,O^{0}[:P_{m}]\}$, representing truncated segments of the full reasoning path at different positions. For each early-exit path, the model generates corresponding answers $C_{1},C_{2},\ldots,C_{m}$.
@@
2.2 Decaying Reward Strategy
To encourage models to produce adequate and correct reasoning steps at earlier stages of CoT generation for accurate problem-solving, we propose a Decaying Reward Strategy. This mechanism assigns rewards based on the correctness of answers generated during the two-time rollouts ( $C_{1},C_{2},\ldots,C_{m},and\ C_{0}$ ), while decaying rewards according to their order of early exits. For the answer $C^{i}$ of each response $O^{i}$, the reward $r^{i}$ is defined as follows:
$r^{i}=\begin{cases}\frac{1}{2^{N_{\text{right}}-1}},&\text{if }C^{i}\text{ is correct,}\\ 0,&\text{if }C^{i}\text{ is incorrect.}\end{cases}$ (1)
Where $N_{\text{right}}$ refers to the accumulated number of correct answers up to and including the current position. The Decaying Reward Strategy is designed with dual objectives: (1) Exponential decay for correct answers: The strategy applies exponentially diminishing rewards to enhance the quality of earlier reasoning steps, which is overlooked by binary 0/1 outcome rewards. (2) Zero reward for incorrect answers: The strategy enforces a correctness-first optimization strategy, ensuring the model maintains robust core reasoning capabilities.
This dual-objective design strikes a balance between reasoning sufficiency and efficiency, guiding the model to produce reasoning sequences that are both accurate and concise.
@@ 2.3 Advantage Computation and Parameter Update
After computing the decaying rewards, GRPO calculates the advantage for each response in the serial group. Specifically, the advantage is calculated by subtracting the mean reward of the group from its corresponding reward, as defined by the formula: $\hat{A}_{i}=r_{i}-\text{mean}(r_{i})$. Here, for training stability, the standard deviation is removed from the advantage computation compared to the GRPO. Subsequently, the computed advantage for each sample is broadcast to all corresponding response tokens. Finally, parameter updates are performed based on the advantage values of each sample (Algorithm 1). The optimization objective is as follows:
@@
5:Sample $m$ temporal positions $P_{i}\in\{P_{1},P_{2},\ldots,P_{m}\}$ where $P_{i}\sim\text{Uniform}(1,n)$
@@
8:for each position $P_{i}$ do
9:     Append $T_{\text{</think>}}$ after position $P_{i}$ to form $CoT^{i}=(T_{1},T_{2},\ldots,T_{P_{i}},T_{\text{</think>}})$
@@ 3.1 Experimental Setup
To comprehensively assess the models’ capabilities across a range of reasoning tasks, we select five popular benchmarks that reflect diverse levels of difficulty: GSM8K (cobbe2021trainingverifierssolvemathgsm8k, ) is a well-curated collection of 1,319 problems in elementary mathematics. This benchmark is specifically designed to evaluate multi-step reasoning in foundational math tasks. Problems typically involve two to eight sequential operations, relying primarily on basic arithmetic performed over multiple intermediate steps. AIME 2024 (aime, ) includes 30 advanced problems selected from the 2024 edition of the American Invitational Mathematics Examination. This highly regarded competition tests participants’ ability to reason mathematically across a broad range of topics, including arithmetic, algebra, combinatorics, geometry, number theory, and probability—core components of secondary-level mathematics. AMC 2023 (AMC2023, ) consists of 40 questions spanning key areas such as algebra, geometry, number theory, and combinatorics. As part of the American Mathematics Competitions organized by the Mathematical Association of America (MAA), AMC aims to foster problem-solving skills and identify promising young mathematicians. MATH-500 math500hendrycks2021measuringmathematicalproblemsolving presents a set of challenging problems drawn from high school-level mathematical competitions. To ensure comparability with prior studies, we use the same subset of 500 problems originally compiled by OpenAI for evaluation purposes (lightman2023letsverifystepstep, ). In addition to these mathematical evaluations, we also examine performance on scientific reasoning tasks, using the following benchmark: GPQA (rein2023gpqagraduatelevelgoogleproofqa, ) is a rigorously constructed dataset containing graduate-level questions in physics, chemistry, and biology. Notably, even domain experts with PhDs achieve only 69.7% accuracy on this benchmark o1 . For our experiments, we focus on the highest-quality subset, GPQA Diamond, which contains 198 carefully chosen questions.
@@
We compare S-GRPO with various existing efficient reasoning methods, including the training-free, output-based approach DEER yang2025dynamicearlyexitreasoning , the off-policy optimization method ConCISE qiao2025conciseconfidenceguidedcompressionstepbystep , and on-policy RL-based approaches such as original GRPO, RL + Length Penalty arora2025traininglanguagemodelsreason , and ShorterBetter yi2025shorterbetterguidingreasoningmodels . DEER makes early-exit decisions during the inference phase, guided by the confidence scores of intermediate answers. ConCISE generates more concise chains of thought by integrating specialized prompt tokens and applying early-exit mechanisms during inference, followed by SFT/SimPO to further encourage succinct reasoning. RL + Length Penalty assigns reward values based on the deviation of each correct response length from the mean, penalizing longer correct responses. ShorterBetter promotes shorter yet accurate reasoning paths by assigning higher rewards to compact chains that yield correct answers based on GRPO.
@@
S-GRPO is designed to improve correctness while minimizing inference length, thereby enabling more efficient reasoning. To evaluate its performance, we adopted two key metrics: Accuracy (i.e., pass@1) and Token Count (Tokens). Due to the inherent instability of generating long sequences in reasoning models and the limited sample sizes of certain benchmarks, we conducted multiple evaluation runs and reported the averaged results in the tables. Specifically, we performed 16 trials on AIME 2024 and AMC 2023, 8 trials on MATH-500 and GPQA Diamond, and 4 trials on GSM8K.
@@
For S-GRPO, we use a learning rate of $1\times 10^{-6}$ and randomly select 8 temporal positions for each query. Since we adopt an on-policy mode, the generation batch size and training batch size are both set to $128\times 8$. For GRPO, we use the same learning rate and batch size settings. For RL + Length Penalty, we follow the settings described in its original paper arora2025traininglanguagemodelsreason and set the scalar parameter $\alpha$ to 0.2. Across all experiments, we employ Adam kingma2014adam as the standard optimizer.
@@
Method GSM8K AIME 2024 AMC 2023 MATH-500 GPQA Overall
@@
Vanilla 92.4 1,833 55.4 13,232 77.2 9,693 85.8 5,590 50.1 15,385 70.02 9,147
DEER 88.8 917 53.3 10,971 87.5 4,142 91.8 2,431 47.5 5,280 73.78+3.76 4,748-48.1%
$\textit{ConCISE}_{SFT}$ 92.9 832 52.1 9,751 92.0 2,244 50.0 5,892
$\textit{ConCISE}_{SimPO}$ 92.1 715 48.3 7,745 91.0 1,946 48.0 4,859
GRPO 93.2 1,767 55.0 13,451 87.5 9,887 93.6 5,317 50.7 15,817 76.00+5.98 9,248+1.1%
RL + Length Penalty 92.4 1,062 51.9 7,464 86.9 3,540 92.2 2,451 49.1 3,984 74.50+4.48 3,700-59.5%
ShorterBetter 53.3 5,288 75.9 2,580
S-GRPO 93.8 906 56.0 7,377 87.5 3,494 92.4 2,252 50.8 3,751 76.10+6.08 3,556-61.1%
@@
Vanilla 94.2 2,129 64.4 11,099 90.5 5,527 93.5 3,844 59.2 6,034 80.36 5,727
DEER 93.3 982 70.0 10,335 90.0 4,349 91.4 2,753 57.1 4,767 80.36+0.0 4,637-19.0%
GRPO 95.3 2,120 65.8 13,504 91.9 6,595 84.0 4,471 58.9 7,354 79.18-1.18 6,809+18.9%
RL + Length Penalty 94.7 775 55.0 7,950 88.1 3,396 92.4 1,993 56.0 4,380 77.24-3.12 3,699-34%
S-GRPO 96.2 724 64.4 6,712 91.9 3,352 93.6 2,146 59.3 3,334 81.08+0.72 3,253-35.4%
@@
Vanilla 95.4 2,370 74.1 15,326 91.3 9,452 93.4 5,577 55.6 8,741 81.90 8,293
DEER 95.5 981 76.7 11,287 95.0 6,198 93.4 3,208 52.5 3,104 82.62+0.72 4,956-40.2%
GRPO 95.8 2,355 72.7 15,154 92.8 8,983 94.4 5,440 55.8 8,819 82.30+0.4 8,150-1.7%
RL + Length Penalty 95.4 1,323 73.8 9,666 93.4 5,042 94.2 3,247 56.2 5,293 82.60+0.7 4,914-40.7%
S-GRPO 96.1 1,292 77.3 8,810 95.0 5,962 95.2 3,166 57.7 5,271 84.26+2.36 4,922-40.6%
@@
Vanilla 95.5 1,909 75.4 14,116 96.9 7,576 95.2 5,078 58.8 7,576 84.36 7,251
DEER 95.5 908 76.7 10,333 95.0 5,099 94.8 2,987 57.1 2,435 83.82-0.54 4,352-40.0%
GRPO 96.1 1,956 77.7 14,544 98.4 8,000 95.8 5,140 59.3 7,966 85.46+1.1 7,521+3.7%
RL + Length Penalty 95.8 1,090 74.8 9,056 96.6 5,059 95.8 2,866 59.4 4,949 84.46+0.1 4,604-36.5%
S-GRPO 96.3 952 77.9 8,932 97.8 4,537 96.4 2,652 60.6 4,537 85.80+1.14 4,322-40.4%